OpenStudy (anonymous):

Find the derivative of f(x) = -12x2 + 9x at x = 6.

OpenStudy (welshfella):

do you know the power rule for finding a derivative?

OpenStudy (welshfella):

if y = ax^n y' = anx^(n-1) this rule can be applied to -12x^2 and + 9x

OpenStudy (welshfella):

for 12x^2 we have a = -12 and n = 2 so to get the derivative of 12x^2 plug these values into the formula for y'

OpenStudy (anonymous):

@welshfella thank you!

OpenStudy (welshfella):

yw

OpenStudy (anonymous):

so to set it up i would do y=-12(2)x^(2-1)?

OpenStudy (anonymous):

where does the 9x come into play?

OpenStudy (anonymous):

@welshfella

OpenStudy (welshfella):

thats right for the derivative of -12x^2 this simplifies to -24x^1 = -24x

OpenStudy (welshfella):

for 9x = 9x^1 a = 9 and n = 1 so derivative is 1*9 x^(1-1) = 9x^0 what is the values of x^0?

OpenStudy (welshfella):

anything to the power 0 = 1 right? so our derivative is 9 so we have f'(x) = -24x + 9